OpenStudy (anonymous):

What is the function inverse of f(x)= 2/3x-5 ?

OpenStudy (anonymous):

y = 2/3x-5 ; find x in terms of y

OpenStudy (anonymous):

x= 2/3y-5? what confuses me is how to rearrange this eqn

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

we need to isolate x. can u take 5 to the other side.

OpenStudy (anonymous):

y(-5)= 2/3x?

OpenStudy (anonymous):

nah, sign will change. y + 5 = 2/3x

OpenStudy (anonymous):

is it it addition? isnt the opposite of division multiplication?

OpenStudy (anonymous):

ya but 5 is subtracted from 2/3x, so it ll be added on the other side rite??

OpenStudy (anonymous):



OpenStudy (anonymous):

ohh sorry. y = 2 / (3x-5) means , 3x - 5 = 2 / y

OpenStudy (anonymous):

(3x-5)/2 = y?

OpenStudy (anonymous):

nah u need to isolate x. 3x = (2 / y) + 5 x = (2 / 3y) + (5/3)

OpenStudy (anonymous):

omg im so confused.

OpenStudy (anonymous):

i thought the 5 was negative

OpenStudy (anonymous):

lemme sum it up for you y = 2 / (3x-5) 3x - 5 = 2 / y 3x = (2 / y) + 5 x = (2 / 3y) + (5/3) Inverse is (2 / 3x) + (5/3)

OpenStudy (anonymous):

okay thanks !

OpenStudy (anonymous):

You're welcome :)